a battery of 10v and negligible internal resistance is coonected across the diagonally opposite corners of a cubical network consisting of 12 resistors each of resistanvce 1ohm. the currnt through the battery is?
 

Grade:12

2 Answers

Arun
5 years ago
Dear Garima
 
Please attach the image as it is required to answer this question.
 
I will be happy to help you.
 
Regards
Arun
Khimraj
5 years ago
efective across the diagonally opposite corners of a cubical network consisting of 12 resistors each of resistanvce 1ohm = 5/6 ohm
Then current through battery = V/R = 10/(5/6) = 12 A
Hope it clears.